### v1

# GEBench: Benchmarking Image Generation Models as GUI Environments

Abstract
Recent advancements in image generation models enable the prediction of future Graphical User Interface (GUI) states based on user instructions. However, existing benchmarks primarily focus on general domain visual fidelity, leaving evaluation of state transitions and temporal coherence in GUI-specific contexts underexplored. To address this gap, we introduce GEBench, a comprehensive benchmark for evaluating dynamic interaction and temporal coherence in GUIs generation. GEBench comprises 700 carefully curated samples spanning five task categories, covering both single-step interactions and multi-step trajectories across real-world and fictional scenarios, as well as grounding point localization. To support systematic evaluation, we propose GE-Score, a five-dimensional metric that assesses Goal Achievement, Interaction Logic, Content Consistency, UI Plausibility, and Visual Quality. Extensive evaluation indicates that current models perform well on single-step transitions but struggle with temporal coherence and spatial grounding over longer interaction sequences. Moreover, our findings identify icon interpretation, text rendering, and localization precision as key bottlenecks, and suggest promising directions for future research toward high-fidelity generative GUI environments. The code is available at: https://github.com/stepfun-ai/GEBench
1 Introduction
Recent advancements in image generation models hurst2024gpt; comanici2025gemini; team2023gemini; seedream2025seedream; flux-2-2025; wan2025 enable the prediction of future Graphical User Interface (GUI) states based on specific user instructions and current visual contexts. This capability positions image generation models as powerful GUI Environments zhang2025large; yan2025gui; xie2025mirage; garg2025controllable; luo2025vimo; wei2023boosting, capable of simulating dynamic interaction sequences to facilitate the scalable training of autonomous agents. Distinguished from conventional simulators cobbe2020leveraging; xie2024osworld; bonatti2024windows tethered to physical hardware or fixed software stacks zhang2025large, these generative models offer a flexible, low-cost alternative for creating diverse interaction trajectories across countless applications zhao2021guigan; liu2025ui.
Figure 1: Comparison of evaluation paradigms across different benchmark types. Existing image generation benchmarks prioritize general-domain visual fidelity and video generation benchmarks evaluate continuous state transitions. GEBench uniquely evaluates discrete state transitions induced by user actions, capturing the essence of GUI interactions.
However, the potential of image generation models as reliable GUI environments remains largely unverified, as traditional visual benchmarks ghosh2023geneval; hu2024ella; niu2025wise; zhao2025envisioning; huang2023t2i; huang2024vbench; sun2025t2v; zhuang2025vistorybench prioritize general-domain visual fidelity (for images) and continuous state transitions (for videos), leaving a critical gap in evaluating the functional logic and state-transition consistency inherent to GUI interactions xie2025gui; yan2025gui. As shown in Figure 1, when acting as GUI environments zhang2025large, generation models must seamlessly navigate discrete, action-triggered interface jumps. Such transitions necessitate precise coordinate grounding zhao2021guigan; cheng2024seeclick, icon recognition liu2025ui; xie2025gui, and high-fidelity text rendering chen2024textdiffuser, compelling the models to maintain logical continuity even when visual elements do not persist li2025mobileworldbench. Such demands strain existing architectures and call for a new evaluation approach to verify if generated GUIs respond felicitously to user instructions.
To bridge this gap, we present GEBench (Benchmarking image generation models as GUI Environments), a benchmark designed to evaluate how effectively image generation models can serve as GUI environments. GEBench comprises 700 high-quality samples, where each entry aligns a sequence of GUI images with corresponding user instructions. These samples span five distinct tasks, allowing for a multifaceted assessment of the model’s ability. To provide a concrete measure of generative quality, we propose GE-Score, a multi-dimensional metric derived from Vision Language Model (VLM)-guided google2025gemini3; hurst2024gpt; bai2025qwen3vltechnicalreport evaluations across five specialized rubrics. GE-Score systematically validates intent fulfillment and interaction logic while verifying UI content consistency and structural integrity. By ensuring high visual fidelity and logical coherence, GE-Score confirms the practical utility of these synthetic environments.
Our systematic evaluation of state-of-the-art image generation models openai2025gptimage; team2023gemini; seedream2025seedream; seedream2025seed; flux-2-2025; wan2025; wu2025qwen; deng2025emerging; li2025uniworld; team2025longcat identifies promising pathways for their evolution into reliable GUI environments. While current architectures demonstrate robust proficiency in executing localized, single-step state transitions, they offer significant opportunities for advancement in long-term interaction consistency and precise spatial grounding. In particular, deficiencies in icon interpretation, Chinese text rendering, and grounding point localization lead to layout drift and logical inconsistencies. These observations delineate critical bottlenecks and outline clear directions for future research toward high-fidelity, temporally coherent generative GUI systems.
Our primary contributions are as follows:
  1. 1.
    We introduce GEBench, a systematic benchmark with 700 samples across five task categories to evaluate image generation models as dynamic GUI environments.
  2. 2.
    We propose GE-Score, a five-dimensional metric that emphasizes the quality assessment of image sequences by accounting for the unique visual properties of GUIs.
  3. 3.
    Our evaluation reveals critical deficiencies in current image generation models, underscoring significant room for improvement in high-fidelity GUI generation.
2 Related Work
2.1 Automated GUIs Generation
The evolution of GUIs generation reflects a significant paradigm shift from heuristic-based structural mapping to data-driven synthesis powered by Multimodal Large Language Models (MLLMs) chen2018ui; sandhaus2011employing; yang2016automatic; li2020layoutgan; zhao2021guigan; mozaffari2022ganspiration; sobolevsky2023guilget; zhang2025scaling; kolthoff2025guide; kolthoff2024zero. Early methodologies relied on traditional rule-based algorithms to perform layout reconstruction sandhaus2011employing; huang2016automaticly, yet these approaches frequently failed to capture the semantic depth of complex hierarchies. Subsequent frameworks simplified this process using model-based approaches to translate visual features directly into code sequences chen2018ui. Contemporary research leverages Transformer-based architectures to bridge the gap between visual design abstractions and executable source code sobolevsky2023guilget; kolthoff2025guide. Furthermore, the rapid advancement of generative AI suggests that direct utilization of image generation models for GUI synthesis is becoming increasingly viable li2020layoutgan; zhao2021guigan; mozaffari2022ganspiration; zhang2025latent. These models offer the potential to produce high-fidelity GUIs directly from user instructions.
2.2 Advanced Image Generation Models
Recent progress in image generation exhibits a rapid evolution from text-to-image synthesis to sophisticated reference-based frameworks hurst2024gpt; comanici2025gemini; team2023gemini; seedream2025seedream; wan2025; deng2025emerging; liu2025step1x; li2025uniworld; team2025nextstep; team2025longcat. Ongoing advancements in text-to-image synthesis have empowered models to produce aesthetically superior visuals with precise semantic alignment to the provided instructions ho2020denoising; chen2020generative; flux-2-2025; fan2024fluid; han2025infinity; ramesh2022hierarchical; lin2025perceiveanythingrecognizeexplain. Building on these foundations, reference-based techniques integrate visual priors with textual prompts to enhance generative control team2023gemini; seedream2025seedream; wan2025; team2025longcat; an2025unictokens. These methods incorporate style or structural references to ensure spatial precision and identity consistency deng2025emerging; liu2025step1x. These advances enable image generation models to function as interactive GUI environments.
2.3 Sequential Generation Benchmarks
Standard image generation benchmarks focus on visual fidelity and text-alignment for a single image, using metrics like FID and CLIP score to measure visual quality heusel2017gans; ghosh2023geneval; huang2023t2i; clip_metrics2021. While these metrics provide a robust baseline for aesthetic realism heusel2017gans, they set the stage for incorporating logical coherence to achieve a more holistic evaluation. Recent efforts in benchmarking sequential image generation explore temporal consistency and reasoning niu2025wise; hu2024ella; zhao2025envisioning; huang2023t2i; zhuang2025vistorybench; zhang2026vibe-benchmark; guo2025video, yet these frameworks typically target natural scenes with continuous movement, simple spatial relationships or characters identity liu2025step1x. GUI environments differ fundamentally because they involve discrete state jumps where a single action replaces the entire visual layout zhang2025large; yan2025gui. Furthermore, the stringent text-rendering requirements of GUIs push current generative architectures to their limits chen2024textdiffuser. This necessitates a new evaluation approach to bridge the significant gap in assessing whether image generation models can maintain the strict semantic and structural integrity required for multi-step GUI trajectories generation.
Figure 2: Examples of the five task types in GEBench, which are designed to comprehensively evaluate the capabilities of image generation models as GUI environments. GEBench provides image generation models with user instructions and reference GUI state (no reference provided for the Fiction App task) and evaluates the generated GUIs.
3 GEBench
A GUI environment functions as an interactive medium that translates user instructions or agent actions into corresponding visual feedback through software logic zhang2025large. This mechanism allows a system to simulate the evolution of a digital interface in response to user interventions. The highly structured and rule-based composition of GUIs, governed by precise functional logic, distinguishes these environments from the patterns of natural scenes.
GEBench establishes an evaluation framework that treats image generation models as interactive GUI environments and benchmarking their performances. Under this paradigm, the model receives visual observations of the current GUI state along with specific user instructions to synthesize the subsequent state. In the following sections, we use terms “GUI state" and “GUIs" to refer to image-based inputs and outputs of image generation models.
3.1 Benchmark Design and Task Suites
GEBench comprises 700 high-quality interaction sequences curated under strict consistency and fidelity constraints. As shown in Figure 2, by organizing samples into five task categories, GEBench enables a fine-grained evaluation of model capabilities across multiple dimensions of GUIs generation:
  1. 1.
    Single-step Visual Transition (single-step) The model receives an initial GUI state as reference image and a detailed action specification to generate the subsequent GUI state. This task evaluates fine-grained instruction following of the models.
  2. 2.
    Multi-step Planning (multi-step) Starting from an initial GUI state as reference and a high-level user objective such as “Order a coffee", the model must generate a five-step GUIs trajectory. This task assesses long-horizon planning, temporal coherence, and the consistency of UI structure across multiple steps.
  3. 3.
    Zero-shot Virtual GUI (fiction-app) This task evaluates out-of-distribution generalization of models, testing whether they can generate unseen layouts by relying on detailed instructions without external reference.
  4. 4.
    Rare Trajectory Synthesis (real-app) This task evaluates the model’s ability to generate long-tail interaction trajectories by prioritizing logical reasoning over pattern imitation, particularly in data-scarce scenarios.
  5. 5.
    Grounding-based Generation (grounding) The model generates the next GUI state based on a normalized relative coordinates within the range of $[0,1000]$. This task assesses spatial awareness and the ability to render changes at precise pixel locations.
Unless otherwise indicated, these five task types are subsequently referred to by their respective parenthetical abbreviations for brevity.
Figure 3: GEBench data construction pipeline. The process involves raw data capture through recording user interactions, task annotation of actions, quality control via preprocessing and verification, and data construction across five task categories: Single-Step, Multi-Step, Grounding, Real App, and Fictional App, totaling 700 samples.
3.2 Evaluation Dimension and Scoring Rubric
GEBench adopts a multi-dimensional scoring rubric designed to assess model abilities in GUIs generation setting. Rather than relying on a single correctness signal, we decompose model performance into five complementary dimensions that jointly capture functional accuracy, interaction realism, and visual fidelity. This design enables fine-grained and interpretable comparisons across all 5 task types.
Specifically, model generated GUI states are evaluated along the following five dimensions:
  • Goal Achievement (GOAL) assesses whether the generated GUI state satisfies the specified action or global objective, focusing on the correctness and completeness of the intended outcome.
  • Interaction Logic (LOGIC) evaluates the plausibility and coherence of state transitions with respect to realistic GUI interaction patterns, ensuring that visual changes can be explained by valid user actions.
  • Consistency (CONS) measures the preservation of unaffected regions within a single image or the stability of UI elements across multiple GUIs, reflecting resistance to unintended visual drift.
  • UI Plausibility (UI) examines whether generated GUI state components are structurally coherent, native-looking, and free from hallucinated or physically impossible elements.
  • Visual Quality (QUAL) evaluates the perceptual fidelity of the generated GUIs, including text readability, icon clarity, and the absence of rendering artifacts.
All dimensions are scored on a discrete ordinal scale from 0 to 5, where higher scores indicate stronger alignment with expected GUIs behavior and visual realism. Detailed scoring criteria for each dimension are provided in the appendix.
We synthesize these multi-dimensional assessments into GE-Score, a holistic metric reflecting the aggregate performance. For each i-th sample evaluated on the d-th dimension, let $r_{i,d}\in\{0,\dots,5\}$ represent the discrete fidelity score. The GE-Score is formally defined as
$GE\penalty 10000\ score=\frac{1}{|\mathcal{D}|\cdot N}\sum_{d\in\mathcal{D}}\sum_{i=1}^{N}\left(\mathcal{F}(r_{i,d})\right)=\frac{4}{N}\sum_{d=1}^{5}\sum_{i=1}^{N}r_{i,d}$ (1)
where $\mathcal{F}(r)=20\times r$ represents the linear normalization transform into percentage domain $[0,100]$. This formulation effectively captures the mean semantic-structural alignment of generation models across the entire benchmark distribution.
3.3 Data Construction Pipeline
The construction of GEBench data follows a structured pipeline designed to transform raw screen recordings into high-quality trajectories for GUI-based interaction. As Figure 3 illustrates, the process begins with the collection of raw interaction data through screen recordings on both mobile and desktop platforms. During the task annotation phase, annotators define specific actions, such as clicking icons or scrolling through interfaces, and convert these sequences into structured JSON metadata.
To further improve data quality, we incorporate a three-stage quality control mechanism. First, a rule-based preprocessing step automatically filters out inconsistent or noisy samples. Second, human experts manually verify the remaining sequences to ensure that the annotated actions accurately match the visual state transitions. Finally, a statistical calibration process adjusts the data distribution to mitigate potential biases, resulting in a final collection of 700 refined samples, which are categorized into five types of tasks, as mentioned in Section 3.1.
Table 1: Main evaluation results on GEBench across Chinese and English Subsets. The table presents a performance comparison across five core dimensions involving 8 commercial models and 4 open-source models. Orange and Champagne cells indicate the Top 1 and Top 2 performers respectively.
Model Chinese Subset English Subset
single multi fiction real ground GE single multi fiction real ground GE
-step -step -app -app -ing Score -step -step -app -app -ing Score
Nano Banana pro team2023gemini 84.50 68.65 65.75 64.35 64.83 69.62 84.32 69.51 46.33 47.20 58.64 61.20
Nano Banana team2023gemini 64.36 34.16 64.82 65.89 54.48 56.74 64.80 50.75 48.88 47.12 49.04 52.12
GPT-image-1.5 openai2025gptimage 83.79 56.97 60.11 55.65 53.33 63.22 80.80 58.87 63.68 58.93 49.23 63.16
GPT-image-1.0 openai2025gptimage 64.72 49.20 57.31 59.04 31.68 52.39 60.92 64.33 58.94 56.16 37.84 55.64
Seedream 4.5 seedream2025seed 63.64 53.11 56.48 53.44 52.90 55.91 49.49 45.30 53.81 51.80 49.63 50.01
Seedream 4.0 seedream2025seedream 62.04 48.64 49.28 50.93 53.53 52.88 53.28 37.57 47.92 49.36 44.17 46.46
Wan 2.6 wan2025 64.20 50.11 52.72 50.40 59.58 55.40 60.17 44.36 49.55 44.80 53.36 50.45
Flux-2-pro flux-2-2025 68.83 55.07 58.13 55.41 50.24 57.54 61.00 52.17 49.92 47.16 45.67 51.18
Bagel deng2025emerging 34.84 13.45 27.36 33.52 35.10 28.85 32.91 8.61 26.08 35.12 37.30 28.00
UniWorld-V2 li2025uniworld 55.33 24.95 32.03 21.39 49.60 36.66 42.68 14.14 30.08 26.83 47.04 32.15
Qwen-Image-Edit wu2025qwen 41.12 26.79 23.78 26.10 50.80 33.72 40.12 18.61 25.80 25.95 54.55 33.01
Longcat-Image team2025longcat 48.76 12.75 30.03 30.00 51.02 34.51 36.69 8.44 37.30 36.83 47.12 33.28
4 Evaluation
4.1 Evaluation Setup
Figure 4: Performance of models across GEBench task suites. The radar chart illustrates the performance of 12 prominent image generation models, including commercial models (solid line) and open-sourced models (dashed line). The reported results represent the average scores on Chinese and English subsets.
Evaluated Models.
Our experimental evaluation encompasses 12 image generation models, grouped into two categories based on accessibility: 8 commercial models and 4 open-source models.
The commercial model group includes Nano Banana pro team2023gemini, Nano Banana team2023gemini, GPT-image-1.5 openai2025gptimage, GPT-image-1 openai2025gptimage, Seedream 4.5 seedream2025seed, Seedream 4.0 seedream2025seedream, Wan2.6 wan2025, Flux-2-pro flux-2-2025.
The open-source model group includes Bagel deng2025emerging, UniWorld-V2 li2025uniworld, Qwen-Image-Edit wu2025qwen, LongCat-Image team2025longcat.
VLM-based Judges.
To ensure the objectivity and robustness of GEBench, we deploy 3 state-of-the-art VLMs as independent cross-evaluators: 2 commercial models Gemini-3-Flash-Native google2025gemini3, GPT-4o hurst2024gpt and 1 open-source model Qwen3-vl-235b-a22b-thinking bai2025qwen3vltechnicalreport. By utilizing these evaluators, we mitigate potential bias inherent in a single judge model. To ensure fair and reproducible comparisons, we use official default configurations for evaluated models and perform evaluation three times for each generated GUIs trajectory.
4.2 Evaluation Results
Overall Performance and Model Comparison.
Experimental results in Table 1 show that Nano Banana Pro team2023gemini delivers the most robust performance, particularly on Chinese subset with a top-ranking GE-Score of 69.62. GPT-image-1.5 openai2025gptimage follows closely, excelling on English subset and securing the first position with a score of 63.16. The radar chart, as shown in Figure 4, further illustrates that commercial models, led by Nano Banana Pro team2023gemini, exhibit a balanced and “full" pentagonal profile. In contrast, open-source models (e.g., UniWorld-V2 li2025uniworld, Bagel deng2025emerging) show performance curves that significantly shrink inward, revealing a substantial gap in handling complex tasks.
The Performance Gap in Multi-step Planning.
The evaluation results in Table 1 highlights a critical bottleneck: while most models excel in Single-step transitions, with both Nano Banana Pro team2023gemini and GPT-image-1.5 openai2025gptimage exceeding 80 points, their scores plummet in Multi-step Planning scenarios, generally dropping below 60 or even 10 points. The radar chart, as shown in Figure 4, clearly identifies the Multi-step axis as a general weakness across nearly all models.
This phenomenon underscores current limitations in long-horizon logical planning:
  • Imbalance between Perception and Planning: Models possess strong instruction-following capabilities for single action mapping but struggle to maintain logical consistency across long sequences.
  • Error Accumulation: During multi-step transitions, minor visual deviations in intermediate steps snowball over time. This accumulated error eventually causes the generated trajectory to diverge entirely from the intended goal.
  • Deficiency in Visual-level Reasoning: Despite their ability to reason within complex textual spaces, models fail to logically grasp inter-step dependencies in interconnected visual tasks, making it difficult to predict the impact of current actions on subsequent visual states.
Challenges in Spatial Grounding.
Cross-metric analysis reveals a general deficit in Grounding-based Generation tasks. As shown in Figure 5, the GOAL (Goal Achievement) score is remarkably low: even top-performing Nano Banana Pro team2023gemini achieves only 23.9%, while most other models (e.g., Qwen-Image-Edit wu2025qwen) fall below 10%. The pronounced “dent" on the Grounding axis of the radar chart, as in Figure 4, further validates this bottleneck: The dismal GOAL scores indicate a logical disconnect; models identify what to generate but cannot translate this into where to place it on a precise $[0,1000]$ coordinate grid. This suggests models lack a fundamental understanding of the mapping between abstract coordinates and image pixels.
Figure 5: Comparison of GOAL score on grounding task. The universally low scores across all models highlight a critical deficiency in current generative models’ ability to perceive and align with precise spatial grounding points.
Figure 6: Pearson correlation analysis between human expert scores and VLM-based evaluation. Results for Nano Banana Pro team2023gemini and GPT-Image-1 openai2025gptimage demonstrate a strong alignment between the VLM-as-a-Judge framework and human judgment across different models.
4.3 Validity of VLM-as-a-Judge
To validate the reliability of using VLM as evaluators, we analyze the correlation between VLM-based assessments and human expert judgments. Specifically, we randomly sample results from two representative models, Nano Banana Pro team2023gemini and GPT-Image-1 openai2025gptimage. For each model, we select 10 edited samples from each of the 10 tasks, resulting in 100 evaluated samples per model. Four human experts independently assess all selected samples using the same evaluation criteria and metrics as the VLM-based judges. Human scores are obtained by averaging all scores across four experts. We then compute the Pearson Correlation Coefficient between the human-annotated scores and the scores produced by the VLM-based judges.
As shown in Figure 6, our VLM-based evaluations exhibit strong correlation with human judgments. The overall Pearson correlation coefficient across all samples reaches r = 0.9892. When analyzed by model, the correlation remains consistently high, with r = 0.9926 for Nano Banana Pro team2023gemini and r = 0.9833 for GPT-Image-1 openai2025gptimage. These results indicate a high level of agreement between the VLM evaluator and human experts across different tasks and models. Together, this analysis demonstrates that the proposed VLM-as-a-Judge framework provides reliable and human-aligned evaluations across diverse tasks and models.
Figure 7: Qualitative results of the three primary weaknesses identified in image generation models acting as GUI environments. The comparison highlights significant deficiencies in text rendering accuracy, icon interpretation for state transitions, and localization precision regarding coordinate-based grounding.
5 Discussion and Analysis
5.1 Hierarchy of Task Difficulty: From Local Mimicry to Global Reasoning Failure
The experimental results, as shown in Table 1 and Figure 4, reveal a pronounced inverse correlation between task complexity and model performance, highlighting a fundamental deficiency in current models’ deep understanding of GUI mechanics. In single-step transition scenarios, leading models team2023gemini; openai2025gptimage achieve robust scores exceeding 80% through powerful visual synthesis. However, this “illusory prosperity" is largely driven by a shortcut mapping from instructions to visual patterns; rather than mastering the underlying interactive logic, models primarily leverage statistical distribution fitting to match expected GUI states. This structural weakness becomes evident as tasks extend to multi-step trajectories. The significant inward “shrinkage" of the radar chart demonstrates an inability to maintain temporal coherence, where the lack of explicit state-space logic leads to a severe logical disconnect when processing high-complexity interaction flows.
5.2 In-depth Bottleneck Analysis: Qualitative Insights from Failure Cases
By synthesizing qualitative evidence, as shown in Figure 7, we identify three primary technical bottlenecks that hinder reliable GUIs generation. The first is the failure of text rendering accuracy. Qualitative cases reveal that while models like Nano Banana Pro team2023gemini encounter deformations in complex layouts, open-source models frequently exhibit severe character overlapping and semantic corruption. This suggests that models treat text as a local texture rather than a symbolic unit with rigid structural information, leading to unreadable characters in layout-dense environments without hard topological constraints.
Secondly, icon interpretation and consistency remain major barriers. Models exhibit significant difficulty in the semanticization of visual symbols, often failing to recognize correct interactive boundaries even when instructions target specific icons. This instability results in “functional distortion" during state transitions, where a specific trigger may degenerate into a meaningless geometric shape. Such a severance of interaction intent and affordance renders interaction entry points unrecognizable for downstream interactions, causing an irreversible break in the task chain.
Finally, a lack of localization precision leads to a critical logical disconnect. Even with explicit coordinate points, generated response elements like pop-up menus exhibit significant spatial jitter, often offsetting by dozens of pixels from their intended locations. This confirms a decoupling of perception and execution, as evidenced by GOAL scores generally falling below 20%. This “blindness" to abstract spatial instructions remains the most formidable obstacle to achieving functionally valid generative GUI environments.
5.3 The Paradox of Visual Fidelity vs. Functional Plausibility
Multi-dimensional analysis through GE-Score reveals a critical paradox: visual fidelity does not equate to functional viability. Models such as GPT-image-1.5 openai2025gptimage generate GUIs with exceptional composition and clarity, earning high QUAL scores. However, granular functional inspection reveals that this “visual over-optimism" is often deceptive, as these aesthetically pleasing images frequently contain hallucinated widgets or illogical layouts. This reinforces that benchmarking image generation models as GUI environments must be predicated upon a assessment of temporal coherence and interactive logic, which take precedence over general-domain visual fidelity.
6 Conclusion
We introduces GEBench, the first systematic benchmark designed to evaluate image generation models as GUI environments. By shifting the focus from general-domain visual fidelity to GUI interaction logic, we provide a comprehensive testbed for assessing the potential of generative models as GUI simulators. Existing image generation models often struggle with the precise structural requirements of interactive GUI generation, a gap that GEBench is uniquely positioned to measure. Through the proposed GE-Score and a VLM-based evaluation pipeline, we identify critical barriers to high-fidelity GUI simulation. Our analysis highlights that while image generation models show promise in predicting basic state transitions, they suffer from icon hallucinations, coordinate drift, and text rendering limitations that hinder their application as robust GUI generators. These findings underscore the need for future research to prioritize fine-grained structural control and semantic persistence over simple visual realism. We believe that GEBench establishes a necessary foundation for developing the next generation of generative GUI simulators capable of supporting the large-scale training of autonomous GUI agents.
Appendix
Appendix A Evaluation Framework
In this section, we detail the proposed evaluation framework. The evaluation framework operates through a systematic three-stage pipeline designed to rigorously benchmark image generation models as GUI environments. The process initiates with Image Generation, where image generation models are tasked with generating visual outputs across five distinct task categories: Single-step Visual Transition (single-step), Multi-step Planning (multi-step), Zero-shot Virtual GUI (fiction-app), Rare Trajectory Synthesis (real-app), and Grounding-based Generation (grounding). Subsequently, the generated samples undergo a VLM-as-a-Judge process employing a evaluator strategy. We leverage VLMs, specifically GPT-4o hurst2024gpt, Gemini-3-Pro-Native google2025gemini3 and Qwen3-vl-235b-a22b-thinking bai2025qwen3vltechnicalreport, to assess the results along five critical dimensions: Goal Achievement (GOAL), Interaction Logic (LOGIC), Consistency (CONS), UI Plausibility (UI), and Visual Quality (QUAL). Finally, the framework concludes with Metrics Analysis, where the calculated scores are validated through statistical verification, pattern analysis, and human relevance alignment to ensure robust and meaningful benchmarking results.
Figure A1: GEBench Evaluation Framework Overview. This diagram outlines the comprehensive evaluation process of the GEBench framework, which assesses the performance of image generation models in generating GUI sequences.
Appendix B Detailed Performance On GEBench Using Different Judges
In this section, we provide a granular breakdown of the experimental results, evaluated by three distinct state-of-the-art Visual Language Models (VLMs) acting as autonomous judges. To ensure the robustness and objectivity of our benchmarking, we report the full performance profiles of all 12 evaluated models across three separate evaluation runs:
  • Table A1: Detailed scores assigned by Gemini-3-Pro-Native google2025gemini3 as the judge.
  • Table A2: Detailed scores assigned by GPT-4o hurst2024gpt as the judge.
  • Table A3: Detailed scores assigned by Qwen3-vl-235b-a22b-thinking bai2025qwen3vltechnicalreport as the judge.
Each table provides a comprehensive matrix of scores across all five GE-Score dimensions (GOAL, LOGIC, CONS, UI, and QUAL) for every task category in GEBench. For consistent comparison and to mitigate the variance in internal scoring scales across different VLM judgers, all raw evaluation outputs have been linearly normalized to a standard range of $[0,100]$. This multi-judger approach allows for a cross-validation of model capabilities and highlights the consistency of our GE-Score framework.
Appendix C Detailed Rubric on five tasks
To ensure a rigorous and standardized evaluation, we developed a series of fine-grained scoring rubrics tailored to the specific requirements of different GUI generation tasks. These rubrics serve as the foundational logic for our VLM-as-a-judge framework.
Each rubric decomposes the five GE-Score dimensions (GOAL, LOGIC, CONS, UI, and QUAL) into explicit, linguistic descriptions across multiple performance tiers. This structured approach minimizes the subjective bias of the VLM judges by providing concrete visual and functional benchmarks for each score level.
Table A1: Detailed Performance on GEBench using Gemini-3-Pro-Native google2025gemini3 as Judge
Task Metric
Nano
Banana Pro
Nano
Banana
GPT-Image
-1.5
GPT-Image
-1.0
Seedream
4.5
Seedream
4.0
Wan 2.6
Flux-2-Pro
Bagel
UniWorld
Qwen
-Image-Edit
Longcat
-Image
single-step GOAL 95.04 69.60 87.65 86.50 56.60 54.58 58.94 65.08 24.35 41.00 25.60 36.07
LOGIC 82.04 58.50 86.50 78.80 49.25 45.29 52.17 53.60 29.12 43.33 27.93 35.47
CONS 75.24 69.70 72.65 58.10 66.23 65.68 81.81 77.46 54.76 58.77 54.93 53.73
UI 79.50 51.60 84.76 58.00 50.19 47.87 53.50 48.43 28.44 47.93 40.07 42.60
QUAL 90.22 73.50 79.97 32.00 61.89 68.39 64.65 79.70 30.95 54.00 52.00 45.73
Overall 84.41 64.58 82.31 62.68 56.83 56.36 62.21 64.85 33.52 49.01 40.11 42.72
multi-step GOAL 86.63 47.61 79.13 77.45 68.25 60.14 46.89 67.18 3.85 2.29 1.98 1.78
LOGIC 56.75 30.46 36.94 35.31 36.63 38.59 25.54 52.71 8.87 3.73 6.70 2.96
CONS 56.95 45.69 36.39 34.08 38.66 38.70 56.43 46.46 15.46 31.86 43.32 19.89
UI 54.96 26.80 50.93 50.41 35.15 33.44 37.56 29.79 12.10 20.17 29.17 16.08
QUAL 90.08 62.03 86.23 85.82 66.80 48.04 64.88 63.18 13.54 40.41 31.60 10.41
Overall 69.07 42.52 57.92 56.61 49.10 43.78 46.26 51.86 10.76 19.69 22.55 10.22
fiction-app GOAL 40.51 40.40 46.20 38.78 40.00 39.33 36.49 45.93 15.80 4.00 0.40 17.01
LOGIC 51.03 48.28 43.13 40.61 39.73 39.80 42.54 48.60 29.80 4.60 0.81 20.90
CONS 57.95 60.00 53.20 53.88 61.00 64.68 61.17 50.67 41.20 88.07 62.36 85.00
UI 50.26 45.45 68.72 61.63 42.60 39.40 41.24 39.00 27.60 28.03 27.95 25.42
QUAL 89.49 89.70 98.20 95.71 92.40 59.87 74.16 85.93 19.20 30.47 32.39 20.90
Overall 57.85 56.77 61.89 58.12 55.15 48.62 51.12 54.03 26.72 31.03 24.78 33.85
real-app GOAL 28.73 32.80 26.96 34.75 40.00 43.80 33.16 34.15 18.20 2.06 0.40 17.07
LOGIC 44.40 45.20 36.36 39.80 37.20 43.27 41.10 47.24 33.80 2.06 0.40 20.00
CONS 67.13 65.33 56.83 56.77 55.00 67.87 59.07 50.71 63.60 72.23 67.67 82.79
UI 48.00 46.80 70.24 58.79 41.20 40.60 39.24 39.25 34.20 22.54 27.60 25.99
QUAL 90.60 92.40 96.16 97.84 89.80 55.20 68.78 84.25 21.80 21.79 32.80 20.88
Overall 55.77 56.51 57.31 57.59 52.64 50.15 48.27 51.12 34.32 24.14 25.77 33.35
grounding GOAL 23.88 21.80 15.19 15.51 16.12 15.42 11.74 11.29 5.86 4.00 4.95 7.47
LOGIC 36.94 41.00 37.39 37.55 51.84 44.58 35.65 36.13 17.17 18.20 14.43 15.35
CONS 77.96 71.20 51.20 45.31 64.29 67.71 83.26 69.68 62.42 69.80 80.21 74.95
UI 77.35 52.00 66.90 61.77 48.57 48.96 68.04 50.32 41.62 70.40 80.41 74.34
QUAL 92.24 72.80 85.57 81.50 75.51 67.08 83.26 79.35 53.94 79.20 83.09 73.13
Overall 61.67 51.76 51.25 48.33 51.27 48.75 56.39 49.35 36.20 48.32 52.62 49.05
Table A2: Detailed Performance on GEBench using GPT-4o hurst2024gpt as Judge
Task Metric
Nano
Banana Pro
Nano
Banana
GPT-Image
-1.5
GPT-Image
-1.0
Seedream
4.5
Seedream
4.0
Wan 2.6
Flux-2-Pro
Bagel
UniWorld
Qwen
-Image-Edit
Longcat
-Image
single-step GOAL 90.73 74.57 88.37 80.41 63.37 60.56 71.38 73.10 31.40 47.30 37.40 40.28
LOGIC 94.25 81.10 92.04 84.39 68.91 66.92 77.18 78.07 38.00 52.20 42.40 44.68
CONS 84.52 79.49 78.27 68.88 70.40 70.32 84.34 83.25 55.90 62.20 64.60 61.44
UI 91.69 78.09 88.57 82.96 66.00 65.55 75.84 78.48 42.20 55.40 55.60 51.00
QUAL 78.95 69.54 78.98 77.14 55.37 58.71 69.68 77.06 37.80 53.20 56.90 48.56
Overall 88.03 76.56 85.25 78.76 64.81 64.41 75.68 77.99 41.06 54.06 51.38 49.19
multi-step GOAL 89.83 72.30 95.44 95.31 88.92 76.96 77.21 83.74 18.31 26.21 15.63 14.01
LOGIC 84.99 72.80 87.16 87.45 80.43 70.14 73.03 78.38 26.83 36.56 32.28 29.53
CONS 83.12 77.70 82.77 82.04 77.70 68.34 81.89 80.51 31.47 56.11 68.36 44.75
UI 89.80 76.60 89.16 89.80 84.39 71.55 80.88 79.49 27.27 52.60 55.82 33.87
QUAL 90.68 78.40 92.32 93.98 85.83 64.68 79.02 76.97 25.37 56.11 42.07 21.31
Overall 87.68 75.56 89.37 89.72 83.45 70.33 78.41 79.82 25.85 45.52 42.83 28.69
fiction-app GOAL 40.36 42.07 51.60 49.21 48.21 40.47 40.14 47.61 23.00 14.53 6.33 25.21
LOGIC 62.28 61.48 70.07 65.15 62.13 55.46 60.89 64.53 43.40 24.87 17.07 46.94
CONS 79.77 76.14 85.33 79.17 79.87 71.20 76.91 72.20 52.60 77.67 68.57 69.58
UI 83.26 79.26 91.28 87.83 81.46 67.72 72.38 72.01 46.87 57.47 45.80 54.20
QUAL 78.10 73.07 87.87 92.32 82.46 52.00 66.80 65.26 35.47 43.40 27.83 40.00
Overall 68.75 66.40 77.23 74.74 70.83 57.37 63.42 64.32 40.27 43.59 33.12 47.19
real-app GOAL 33.40 39.24 43.73 43.83 43.73 41.34 41.42 41.26 22.80 9.11 4.27 25.00
LOGIC 55.80 59.32 61.72 59.06 57.20 55.47 50.25 62.87 50.67 19.42 17.53 43.80
CONS 79.00 80.48 81.66 81.22 76.33 69.67 76.75 76.28 64.13 71.10 69.27 70.80
UI 81.60 81.56 88.40 87.35 78.00 64.93 71.75 76.54 50.73 46.77 42.33 58.80
QUAL 78.40 78.77 88.15 92.40 80.33 48.99 61.92 69.67 34.53 32.85 22.47 41.80
Overall 65.64 67.87 72.73 72.77 67.12 56.08 60.42 65.32 44.57 35.85 31.17 48.04
grounding GOAL 35.60 26.07 20.40 16.33 21.55 17.73 15.85 17.40 8.47 7.89 6.73 5.82
LOGIC 67.00 66.13 53.80 55.80 58.38 58.14 57.21 58.87 54.73 61.00 63.87 62.07
CONS 78.20 85.00 75.82 59.33 75.35 68.04 84.69 79.80 67.07 86.45 91.53 89.03
UI 81.40 77.67 76.03 67.73 69.23 62.89 77.41 71.93 55.13 85.45 91.47 88.23
QUAL 80.40 76.53 80.74 75.93 75.76 63.30 82.18 77.53 55.27 87.17 90.33 84.01
Overall 68.52 66.28 61.36 55.02 60.05 54.02 63.47 61.11 48.13 65.59 68.79 65.83
Table A3: Detailed Performance on GEBench using Qwen3-vl-235b-a22b-thinking bai2025qwen3vltechnicalreport as Judge
Task Metric
Nano
Banana Pro
Nano
Banana
GPT-Image
-1.5
GPT-Image
-1.0
Seedream
4.5
Seedream
4.0
Wan 2.6
Flux-2-Pro
Bagel
UniWorld
Qwen
-Image-Edit
Longcat
-Image
single-step GOAL 93.64 70.54 84.70 75.53 67.24 65.85 67.65 66.84 27.14 38.72 33.23 35.05
LOGIC 92.45 72.86 81.27 71.20 66.02 64.41 69.13 68.27 30.44 38.32 35.32 34.78
CONS 83.28 66.40 64.60 55.43 59.66 57.68 70.31 69.97 37.58 36.50 44.07 34.88
UI 94.25 65.62 83.98 70.14 62.41 61.16 65.02 63.57 31.85 40.17 44.78 35.59
QUAL 98.04 80.51 94.85 86.80 72.59 76.03 77.37 80.58 41.21 55.19 60.30 48.48
Overall 92.33 71.19 81.88 71.82 65.58 65.03 69.90 69.85 33.64 41.78 43.54 37.76
multi-step GOAL 92.35 60.97 93.61 96.22 89.39 83.45 72.70 83.80 4.62 6.44 4.22 3.70
LOGIC 88.53 60.44 84.43 89.46 77.59 75.09 65.92 76.15 17.65 23.95 19.70 16.69
CONS 84.04 71.32 75.50 79.22 73.23 71.26 76.32 72.29 19.69 42.06 57.16 30.29
UI 95.96 60.77 90.68 92.31 75.00 73.23 74.48 64.74 16.12 33.32 51.10 21.16
QUAL 97.65 71.89 96.34 98.20 86.87 80.17 87.13 77.29 17.65 48.06 53.59 21.16
Overall 91.71 65.08 88.11 91.08 80.42 76.64 75.31 74.85 15.15 30.77 37.15 18.60
fiction-app GOAL 28.13 34.48 43.33 34.68 39.19 34.27 27.50 34.13 15.69 9.73 8.07 18.40
LOGIC 41.93 49.90 58.33 46.87 48.26 45.67 45.28 48.70 30.24 19.60 13.93 31.39
CONS 49.47 58.05 52.87 43.30 49.26 43.73 49.51 43.33 25.86 53.07 20.53 36.04
UI 53.47 55.69 70.20 63.91 55.64 45.27 47.64 40.67 22.36 34.00 19.47 29.93
QUAL 76.53 72.73 97.00 94.75 82.21 60.47 73.12 63.00 20.88 34.00 24.47 29.03
Overall 49.91 54.17 64.35 56.70 54.91 45.88 48.61 45.97 23.01 30.08 17.29 28.96
real-app GOAL 30.47 30.47 34.73 37.33 36.67 39.00 27.95 35.08 17.58 8.40 4.33 21.73
LOGIC 43.47 47.32 50.73 49.40 45.87 46.27 47.03 48.56 33.60 14.95 12.80 31.33
CONS 58.33 62.21 53.13 49.93 48.73 42.80 51.30 47.81 34.75 37.75 17.60 43.96
UI 59.80 59.53 73.07 65.47 50.47 41.53 50.88 48.42 26.06 22.39 15.13 34.93
QUAL 75.60 78.52 97.93 93.67 80.13 56.00 67.03 66.94 20.88 22.73 16.60 34.53
Overall 53.53 55.61 61.92 59.16 52.37 45.12 48.84 49.36 26.57 21.24 13.29 33.30
grounding GOAL 24.38 20.89 15.68 18.72 19.31 14.41 16.05 16.58 12.44 12.50 6.69 9.08
LOGIC 36.50 36.56 32.26 30.54 37.66 32.21 34.45 31.81 22.34 22.85 15.00 18.16
CONS 46.20 42.89 33.77 25.10 40.96 38.76 48.21 46.44 35.79 40.83 42.04 42.13
UI 69.49 58.11 56.85 44.30 56.22 52.55 67.22 56.98 40.07 59.03 68.66 61.35
QUAL 89.12 79.38 81.10 70.40 76.29 72.76 86.08 73.56 53.11 81.39 84.23 80.21
Overall 53.14 47.57 43.93 37.81 46.09 42.14 50.40 45.07 32.75 43.32 43.32 42.19
Appendix D Detailed Rubric on five tasks
To ensure a rigorous and standardized evaluation, we developed a series of fine-grained scoring rubrics tailored to the specific requirements of different GUI generation tasks. These rubrics, detailed in Figures A2, A3, A4, and A5, serve as the foundational logic for our VLM-as-a-judge framework.
Each rubric decomposes the five GE-Score dimensions (GOAL, LOGIC, CONS, UI, and QUAL) into explicit, linguistic descriptions across multiple performance tiers (e.g., from “Incomplete" to “Exceptional"). This structured approach minimizes the subjective bias of the VLM judges by providing concrete visual and functional benchmarks for each score level. Specifically:
  • Figure A2 outlines the criteria for Single-Step Transition, focusing on the immediate visual mapping of user instructions.
  • Figure A3 details the Multi-Step Planning rubrics, emphasizing the accumulation of errors and temporal coherence across 5 step trajectories.
  • Figure A4 provides the standards for Zero-shot Virtual GUI generation, where the judge assesses the imaginative plausibility and structural integrity of non-existent applications.
  • Figure A5 defines the benchmarks for Grounded Generation, specifically evaluating the pixel-level alignment between generated content and coordinate-based prompts.
By employing these detailed rubrics, we bridge the gap between qualitative visual inspection and quantitative performance metrics.
Figure A2: Evaluation Rubrics for Single-Step Transition Generation
Figure A3: Evaluation Rubrics for Multi-Step Planning Generation
Figure A4: Evaluation Rubrics for Zero-shot Virtual GUI Generation and Rare Trajectory Synthesis
Figure A5: Evaluation Rubrics for Grounding-based Generation